OpenStudy (anonymous):

Write the first ten terms of a sequence whose first term is -10 and whose common difference is -2.

OpenStudy (texaschic101):

common difference means it is an arithmetic sequence. first term = -10 second term = -10 + (-2) = -10 - 2 = -12 third term = -12 + (-2) = -12 - 2 = -14 and so on...do you understand how I am arriving at the answers ?